# Data-Efficient Fine-Tuning of Vision-Language Models for Diagnosis of Alzheimer’s Disease

### v1

@@ Abstract
Medical vision-language models (Med-VLMs) have shown impressive results in tasks such as report generation and visual question answering, but they still face several limitations. Most notably, they underutilize patient metadata and lack integration of clinical diagnostic knowledge. Moreover, most existing models are typically trained from scratch or fine-tuned on large-scale 2D image-text pairs, requiring extensive computational resources, and their effectiveness on 3D medical imaging is often limited due to the absence of structural information. To address these gaps, we propose a data-efficient fine-tuning pipeline to adapt 3D CT-based Med-VLMs for 3D MRI and demonstrate its application in Alzheimer’s disease (AD) diagnosis. Our system introduces two key innovations. First, we convert structured metadata into synthetic reports, enriching textual input for improved image-text alignment. Second, we add an auxiliary token trained to predict the mini-mental state examination (MMSE) score, a widely used clinical measure of cognitive function that correlates with AD severity. This provides additional supervision for fine-tuning. Applying lightweight prompt tuning to both image and text modalities, our approach achieves state-of-the-art performance on two AD datasets using 1,500 training images, outperforming existing methods fine-tuned on 10,000 images. Code will be released upon publication.
@@ 1 Introduction
Due to the significant gap between natural images and medical images, many Med-VLMs [8, 9, 10, 11] are pre-trained on large-scale datasets consisting of 2D medical images and text from open access papers [12] or medical reports [13]. However, these models often require further adaptation for specific clinical applications. Currently, fine-tuning Med-VLMs for disease-specific tasks generally involves training both image and text encoders [14, 15], which requires substantial amounts of data and computing resources. Additionally, for conditions such as Alzheimer’s disease (AD), which rely on 3D MRI scans, existing models fail to fully exploit the spatial structure inherent in volumetric data. Therefore, developing a data-efficient fine-tuning approach based on pre-trained 3D-Med-VLMs, such as those proposed in [16, 17], is crucial for accurate and clinically practical diagnosis of these diseases.
@@
In this paper, we propose a data-efficient pipeline for fine-tuning 3D Med-VLMs, models originally trained on a variety of medical imaging modalities, for the diagnosis of AD using 3D MRI data. A key challenge in AD diagnosis is the lack of detailed textual reports. Given that the dataset contains metadata describing subject-specific information and MRI-derived attributes, such as hippocampal volume, ventricular size and whole-brain volume, we propose leveraging this information to generate synthetic reports, thereby providing richer textual supervision. Moreover, most existing models lack integration of clinical diagnostic knowledge, which is essential for informed decision-making in real-world settings. Since cognitive assessments, such as mini-mental state examination (MMSE) scores, play a crucial role in clinical diagnosis of AD, we introduce a learnable auxiliary token for predicting the MMSE score. This auxiliary prediction task further guides the tuning of image and text representations. To address the domain gap arising from differences in imaging principles and characteristics between MRI and other radiological modalities, we adopt prompt tuning [18], a lightweight and parameter-efficient fine-tuning approach that adapts both the image and text encoders. Additionally, a cross-attention mechanism is applied to enhance alignment between visual and textual modalities.
@@
We evaluate the proposed fine-tuning pipeline on two AD datasets: ADNI [19] and OASIS [20], focusing on distinguishing AD, mild cognitive impairment (MCI), and normal controls (NC). Using only 1,504 MRIs for fine-tuning, our approach consistently improves classification performance across various pre-trained 3D Med-VLMs. For example, on the ADNI dataset, compared to directly using the backbone models for AD classification, our pipeline improves the accuracy of RadFM by 24.5% and M3D by 23.9%. Even for Med3DVLM, whose encoder was fully fine-tuned on our dataset as a baseline, our method achieves 4.8% higher accuracy. Despite using significantly fewer training samples, our method achieves 78.8% accuracy in this three-way classification task, outperforming a model specifically designed for AD diagnosis and originally trained on 10,387 MRIs. Furthermore, our method outperforms other fine-tuning approaches under the same training conditions. Additionally, visualization results validate that textual input effectively guides the model to focus on clinically meaningful features, such as ventricular size.
@@ 3 Method
Figure 1 illustrates the proposed data-efficient fine-tuning pipeline. For the visual modality, we utilize the pre-trained 3D ViT as the vision encoder to process 3D medical images and incorporate visual prompt learning [18] to adapt from other medical image types (e.g., CT or X-ray) to MRI. An additional MMSE token is included alongside the typical patch embeddings and class token, which is used to predict the MMSE score, mimicking real-life clinical diagnosis and aiding the fine-tuning process. For the text modality, we leverage the transformer-based text encoder [1, 16] to process synthetic reports generated from metadata and introduce learnable prompts to enhance its understanding of AD-related concepts. Prior to calculating image-text similarity, cross-attention is used to dynamically adjust the importance of attributes and improve alignment between the two modalities.
@@ 3.1 Vision encoder
where the patch embedding module, $\mathit{PatchEmbed}$, consists of a 3D convolutional layer with kernel size and stride equal to the patch size, ensuring that each convolution extracts exactly one patch and maps it to an embedding vector, followed by a learnable positional embedding to preserve spatial information. The set of patch embeddings is represented as $\mathbf{E}_{0}=\{e_{0}^{n}\in\mathbb{R}^{d}\mid n=1,2,\ldots,m\}$. These patch embeddings are concatenated with a learnable class token, $\mathit{CLS}]$, and fed into the ViT. The ViT consists of $n_{\ell}$ layers, where the $\ell$-th transformer layer is denoted as $L_{\ell}$.
In addition to the class token, we introduce an auxiliary learnable token, $\mathit{MMSE}]$, specifically trained to predict the mini-mental state examination (MMSE) score. MMSE is a widely used cognitive assessment for AD diagnosis, with lower scores indicating more severe cognitive impairment. Incorporating this prediction task provides additional supervision during training, encouraging the model to extract clinically relevant features from MRI scans and thereby enhancing both performance and interpretability.
@@
$=L_{\ell}([C_{\ell-1},M_{\ell-1},\mathbf{P}_{\ell-1},\mathbf{E}_{\ell-1}]),$ (3)
$\quad\ell=1,2,\ldots,n_{\ell}.$
Where $C_{\ell}$ and $M_{\ell}$ denote the embedding of the $\mathit{CLS}]$ and $\mathit{MMSE}]$ tokens respectively at layer $\ell$, $E_{\ell}$ denotes the patch embeddings. After the final transformer layer, we extract the class and MMSE token embeddings $C_{n_{\ell}}]$ and $M_{n_{\ell}}]$. The class token embedding is then projected into a joint vision-language embedding space via a linear layer followed by $l_{2}$ normalization, denotes as $\textit{f}_{v}$:
@@
For the final MMSE embedding $M_{n_{\ell}}$, we use a prediction head $\textit{f}_{\mathit{mmse}}$ composed of two linear layers with a non-linear activation in between to predict the MMSE score $\hat{y}_{\mathit{mmse}}$:
@@ 3.2 Text encoder
To address the issue of missing MRI-related reports, we generate the report, denoted as $T$, from the MRI’s metadata. The report $T$ is created using attributes extracted from the metadata that reflect changes in brain structure during the progression of the disease. For example, the generated report might read: “A photo of NC. The MRI scan reveals the following biomarkers: Hippocampal volume: 7323.00 $mm^{3}$, Ventricular size: 43767.00 $mm^{3}$, Whole brain volume: 968731.00 $mm^{3}$, Entorhinal cortex volume: 4056.00 $mm^{3}$, Fusiform gyrus volume: 18775.00 $mm^{3}$, Middle temporal gyrus volume: 17048.00 $mm^{3}$.” This kind of text are expected to guide the learning of the image representations to incorporate more accurate information to help the diagnosis of AD.
During processing, $T$ is tokenized into a sequence of tokens using a subword tokenizer. Since the text encoder is pre-trained on natural language and may lack sufficient knowledge of AD, we incorporate the text prompt $\mathbf{R}$ at the input layer to enhance the text encoder $E_{text}$’s understanding of AD-related information. $l_{r}$ is the length of text prompts.
$\mathbf{R}=\{R^{q}_{l}\mid l\in\mathbb{N},q=1,2,3,\ldots,l_{r}\}$ (6)
@@
$\mathbf{w}=\textit{f}_{t}~(E_{text~}~([\mathbf{R},Tokens]))$ (7)
@@ 3.3 Attention-based aggregation
To differentiate the significance of variables in the generated reports, we incorporate cross-attention layers following the projection layers for both image and text embeddings. This allows image features to integrate textual information while refining text features with visual information,dynamically adjusting the attributes importance, for better alignment between MRI and reports. Following [32], we introduce a cross-attention layer to refine the features:
$\hat{\mathbf{v}}=\text{softmax}\left(\frac{\mathbf{v}\cdot\mathbf{w}^{T}}{\sqrt{d_{k}}}\right)\mathbf{w}+\mathbf{v}$ (8)
$\hat{\mathbf{w}}=\text{softmax}\left(\frac{\mathbf{w}\cdot\mathbf{v}^{T}}{\sqrt{d_{k}}}\right)\mathbf{v}+\mathbf{w}$ (9)
@@ 3.4 Loss functions
We expect that image $i$ have the highest similarity with its corresponding text, leading to the image-to-text contrastive loss:
$\mathcal{L}_{\text{image-to-text}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp s_{ii}/\tau}{\sum_{j=1}^{N}\exp s_{ij}/\tau},$ (11)
where $\tau$ is a learnable temperature parameter that controls the sharpness of the similarity distribution. Similarly, we compute the text-to-image loss:
$\mathcal{L}_{\text{text-to-image}}=-\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp s_{jj}/\tau}{\sum_{i=1}^{N}\exp s_{ij}/\tau}.$ (12)
Combining these two loss functions, we get the contrastive learning loss:
$\mathcal{L}_{CL}=\mathcal{L}_{\text{image-to-text}}+\mathcal{L}_{\text{text-to-image}}.$ (13)
@@
$\mathcal{L}=\mathcal{L}_{CL}+\lambda\mathcal{L}_{\mathit{mmse}},$ (15)
where $\lambda$ is a weighting coefficient that balances the relative importance of the contrastive loss $\mathcal{L}_{CL}$ and the MMSE prediction loss $\mathcal{L}\mathit{mmse}$.
@@ 4.1.1 Datasets
During evaluation, we assess all methods on two AD datasets. First, we evaluate on the ADNI dataset using 268 MRI scans from 79 subjects who are not included in fine-tuning, comprising 21 NC, 33 MCI, and 25 AD cases. Second, we evaluate on a subset of the Open Access Series of Imaging Studies (OASIS-2) dataset [20], which contains 160 MRIs from 60 subjects (30 AD, 30 NC). This setup allows for assessing the model’s zero-shot generalizability on an unseen dataset, albeit for a binary classification task.
@@ 4.1.3 Baseline methods
Methods ACC (%) # Parameters (Learnable) # Training samples
ADNI OASIS
MedCLIP [8] 39.6 40.0 200M (0) -
BioMedCLIP [10] 40.7 41.3 210M (0) -
RadFM [21] 52.9 60.4 120M (0) -
   + 3D Spatial Pooling Perceiver ++ LoRA [16] 69.8 68.1 + 35.7M 1,504
   + LGA [34] 74.6 76.3 + 85.2M 1,504
[…]
M3D [16] 54.9 52.6 198M (0)
   + 3D Spatial Pooling Perceiver ++ LoRA [16] 74.3 78.8 + 35.7M 1,504
   + LGA [34] 76.7 80.0 + 85.2M 1,504
[…]
Med3DVLM [25] 71.5 75.2 199M (199M) 1,504
   + 3D Spatial Pooling Perceiver ++ LoRA [16] 74.6 79.4 + 35.7M 1,504
   + LGA [34] 75.0 78.1 + 85.2M 1,504
[…]
MedBLIP [35] 78.7 85.3 1.7B (154M) 10,387
Table 2: Classification accuracy (%) of different backbones with and without our fine-tuning pipeline on ADNI and OASIS datasets. Results are averaged over five runs.
@@ 4.2 Quantitative result of classification task
Finally, we evaluate our method’s zero-shot performance on the OASIS-2 dataset. Our model achieves strong results, demonstrating good generalizability to unseen data without additional fine-tuning. It is important to note that MedBLIP’s original reported results on OASIS are not zero-shot, as it was trained on the OASIS dataset itself, which makes direct comparison less fair.

### v3

@@ Abstract
Medical vision-language models (Med-VLMs) have shown impressive results in tasks such as report generation and visual question answering, but they still face several limitations. Most notably, they underutilize patient metadata and lack integration of clinical diagnostic knowledge. Moreover, most existing models are typically trained from scratch or fine-tuned on large-scale 2D image-text pairs, requiring extensive computational resources, and their effectiveness on 3D medical imaging is often limited due to the absence of structural information. To address these gaps, we propose a data-efficient fine-tuning pipeline to adapt 3D CT-based Med-VLMs for 3D MRI and demonstrate its application in Alzheimer’s disease (AD) diagnosis. Our system introduces two key innovations. First, we convert structured metadata into synthetic reports, enriching textual input for improved image-text alignment. Second, we add an auxiliary token trained to predict the mini-mental state examination (MMSE) score, a widely used clinical measure of cognitive function that correlates with AD severity. This provides additional supervision for fine-tuning. Applying lightweight prompt tuning to both image and text modalities, our approach achieves state-of-the-art performance on ADNI with only 1,504 training MRIs, outperforming methods trained on 27,161 MRIs, and shows strong zero-shot generalization on OASIS-2 and AIBL. Code is available at https://github.com/CFQ666312/DEFT-VLM-AD.
@@ 1 Introduction
Due to the significant gap between natural images and medical images, many Med-VLMs [8, 9, 10, 11] are pre-trained on large-scale datasets consisting of 2D medical images and text from open access papers [12] or medical reports [13]. However, these models often require further adaptation for specific clinical applications. Currently, fine-tuning Med-VLMs for disease-specific tasks generally involves training both image and text encoders [14, 15], which requires substantial amounts of data and computing resources. Additionally, for conditions such as Alzheimer’s disease (AD), which rely on 3D MRI scans, existing models fail to fully exploit the spatial structure inherent in volumetric data. Therefore, developing a data-efficient fine-tuning approach based on pre-trained 3D Med-VLMs, such as those proposed in [16, 17], is crucial for accurate and clinically practical diagnosis of these diseases.
@@
In this paper, we propose a data-efficient pipeline for fine-tuning 3D Med-VLMs to diagnose AD using 3D MRI data. A key challenge in AD diagnosis is the lack of detailed radiology reports. Given that the dataset contains metadata describing subject-specific information and MRI-derived attributes, we propose leveraging this information to generate synthetic reports, thereby providing richer textual supervision. Moreover, most existing models lack integration of clinical diagnostic knowledge, which is vital for informed decision-making in real-world settings. Since cognitive assessments, such as mini-mental state examination (MMSE) scores, play a crucial role in clinical diagnosis of AD, we introduce a learnable auxiliary token for predicting the MMSE score. This auxiliary prediction task further guides the tuning of image and text representations. To address the domain gap arising from differences in imaging principles and characteristics between MRI and other radiological modalities, we adopt prompt tuning [18], a lightweight and parameter-efficient fine-tuning approach that adapts both the image and text encoders. Additionally, a cross-attention mechanism is applied to enhance alignment between visual and textual modalities.
@@
We evaluate the proposed fine-tuning pipeline on three AD datasets, distinguishing AD, mild cognitive impairment (MCI), and normal controls (NC). Using only 1,504 MRIs for fine-tuning, our approach consistently improves classification performance across various pre-trained 3D Med-VLMs. On the ADNI dataset, our pipeline improves RadFM [19] by 24.5%, M3D by 23.9% and Med3DVLM by 25.9%. Notably, it achieves 78.8% accuracy in this three-way classification task, outperforming MedBLIP [20] trained on 27,161 MRIs. Furthermore, our method outperforms other fine-tuning approaches under the same training conditions. Additionally, visualization results validate that textual input effectively guides the model to focus on clinically meaningful features, such as ventricular size.
@@ 3 Method
Figure 1 illustrates the proposed data-efficient fine-tuning pipeline. For the visual modality, we utilize the pre-trained 3D ViT as the vision encoder to process 3D medical images and incorporate visual prompt learning [18] for adapting from other modalities to MRI. An additional MMSE token is included alongside the standard patch embeddings and class token; it predicts the MMSE score, mimicking clinical diagnosis and guiding fine-tuning. For the text modality, we leverage the transformer-based text encoder [1, 16] to process synthetic reports generated from metadata and introduce learnable prompts to enhance its understanding of AD-related concepts. Prior to calculating image-text similarity, cross-attention is used to dynamically adjust the importance of attributes and improve alignment between the two modalities.
@@ 3.1 Vision encoder
where the patch embedding module, $\mathit{PatchEmbed}$, consists of a 3D convolutional layer with kernel size and stride equal to the patch size, followed by a learnable positional embedding. The set of patch embeddings is represented as $\mathbf{E}_{0}=\{e_{0}^{n}\in\mathbb{R}^{d}\mid n=1,2,\ldots,m\}$. These patch embeddings are concatenated with a learnable class token, $\mathit{CLS}]$, and fed into the ViT. The ViT consists of $n_{\ell}$ layers, where the $\ell$-th transformer layer is denoted as $L_{\ell}$.
In addition to the class token, we introduce an auxiliary learnable token, $\mathit{MMSE}]$, trained to predict MMSE score, where lower scores indicate more severe cognitive impairment. This auxiliary prediction provides additional supervision, encouraging the model to extract clinically relevant features from MRI scans.
@@
$C_{\ell},M_{\ell},\__{,}\mathbf{E}_{\ell}]=L_{\ell}([C_{\ell-1},M_{\ell-1},\mathbf{P}_{\ell-1},\mathbf{E}_{\ell-1}]),\\ \quad\ell=1,2,\ldots,n_{\ell},$ (3)
where $C_{\ell}$ and $M_{\ell}$ denote the embedding of the $\mathit{CLS}]$ and $\mathit{MMSE}]$ tokens respectively at layer $\ell$, $\mathbf{E}_{\ell}$ denotes the patch embeddings. After the final transformer layer, we extract the class and MMSE token embeddings $C_{n_{\ell}}]$ and $M_{n_{\ell}}]$. The class token embedding is then projected into a joint vision-language embedding space via a linear layer followed by $l_{2}$ normalization, denotes as $\textit{f}_{v}$:
@@
For the final MMSE embedding $M_{n_{\ell}}$, we use a prediction head $\textit{f}_{\mathit{mmse}}$ composed of two linear layers with a non-linear activation between them to predict the MMSE score $\hat{y}_{\mathit{mmse}}$:
@@ 3.2 Text encoder
A considerable amount of work focuses on radiology report generation, where models learn to produce clinical text from medical images or other large language models [pellegrini2023radialog, gu2025radalign]. These methods typically optimize language generation objectives to produce clinical reports. However, while they can indirectly improve image representations, they are not explicitly designed to extract discriminative features for diagnosis. In contrast, our approach aims to address the issue of missing MRI-related reports and to guide image representation learning for AD, without producing full reports in natural language. Therefore, we generate structured textual descriptions from MRI metadata. The report, denoted as $T$, is created using attributes extracted from the metadata that reflect changes in brain structure during the progression of the disease. For example, the generated report might read: “A photo of NC. The MRI scan reveals the following biomarkers: Hippocampal volume: 7323.00 $mm^{3}$, Ventricular size: 43767.00 $mm^{3}$, Whole brain volume: 968731.00 $mm^{3}$, Entorhinal cortex volume: 4056.00 $mm^{3}$, Fusiform gyrus volume: 18775.00 $mm^{3}$, Middle temporal gyrus volume: 17048.00 $mm^{3}$.”
During processing, $T$ is tokenized into a sequence of tokens, $\mathit{Tokens}$, using a subword tokenizer. Since the text encoder is pre-trained on natural language and may lack sufficient knowledge of AD, we incorporate the text prompt $\mathbf{R}$ of length $l_{r}$ at the input layer to enhance the text encoder $E_{text}$’s understanding of AD-related information:
$\mathbf{R}=\{R^{q}\mid q=1,2,3,\ldots,l_{r}\}.$ (6)
@@
$\mathbf{w}=\textit{f}_{t}~(E_{text~}~([\mathbf{R},\mathit{Tokens}])),$ (7)
@@ 3.3 Attention-based aggregation
To differentiate the significance of variables in the generated reports, we incorporate cross-attention layers following the projection layers for both image and text embeddings. This allows image features to integrate textual information while refining text features with visual information, dynamically adjusting the attributes importance for better alignment between MRI and reports. Following [31], we introduce a cross-attention layer to refine the features:
$\hat{\mathbf{v}}=\text{softmax}\left(\frac{\mathbf{v}\cdot\mathbf{w}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{w}+\mathbf{v},\ \hat{\mathbf{w}}=\text{softmax}\left(\frac{\mathbf{w}\cdot\mathbf{v}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{v}+\mathbf{w},$ (8)
@@ 3.4 Loss functions
We expect image $i$ to have the highest similarity with its corresponding text, and likewise, text $j$ to have the highest similarity with its corresponding image, leading to the image-to-text (I2T) and text-to-image (T2I) contrastive losses:
$\small\mathcal{L}_{\text{I2T}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp s_{ii}/\tau}{\sum_{j=1}^{N}\exp s_{ij}/\tau},\ \mathcal{L}_{\text{T2I}}=-\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp s_{jj}/\tau}{\sum_{i=1}^{N}\exp s_{ij}/\tau},$ (10)
where $\tau$ is a learnable temperature parameter that controls the sharpness of the similarity distribution.
@@
$\mathcal{L}=\mathcal{L}_{\text{I2T}}+\mathcal{L}_{\text{T2I}}+\lambda\mathcal{L}_{\mathit{mmse}},$ (12)
where $\lambda$ is a weighting coefficient that balances the relative importance of the contrastive loss and the MMSE prediction loss.
@@ 4.1.1 Datasets
During evaluation, we assess all methods on three AD datasets. First, we evaluate on the ADNI dataset using 268 MRI scans from 79 subjects who are not included in fine-tuning, comprising 21 NC, 33 MCI, and 25 AD. Second, to assess the model’s zero-shot generalizability on an unseen dataset, we evaluate on a balanced subset of the OASIS-2 dataset [34], which contains 160 MRIs from 60 subjects (30 demented, 30 non-demented), and on a subset of the AIBL dataset [35], which contains 150 MRIs (50 NC, 50 MCI, and 50 dementia).
@@ 4.2 Quantitative result of classification task
Table 2: Classification accuracy of different backbones with and without fine-tuning on the ADNI, OASIS and AIBL datasets, as well as computational cost. Our results are averaged over five random runs and reported as mean $\pm$ standard deviation. $\dagger$ indicates results taken from the original paper.
Methods ACC (%) # Parameters Time Mem.
ADNI OASIS AIBL (Learnable) /Epoch /Card
MedCLIP [8] 39.6 40.0 34.7 200M (0) - -
BioMedCLIP [10] 40.7 41.3 25.3 210M (0) - -
RadFM [19] 52.9 60.4 35.3 120M (0) - -
+ 3D SP Perceiver [16] 69.8 68.1 51.3 + 35.7M 508.6s 43.4G
+ LGA [36] 74.6 76.3 57.3 + 85.2M 932.4s 46.1G
[…]
M3D [16] 54.9 52.6 34.0 198M (0) - -
+ 3D SP Perceiver [16] 74.3 78.8 54.7 + 35.7M 508.6s 43.4G
+ LGA [36] 76.7 80.0 52.0 + 85.2M 932.4s 46.1G
[…]
Med3DVLM [24] 50.4 48.3 44.7 199M (0) - -
+ 3D SP Perceive [16] 74.6 79.4 68.0 + 35.7M 508.6s 43.4G
+ LGA [36] 75.0 78.1 67.3 + 85.2M 932.4s 46.1G
[…]
MedBLIP [20] 78.7 85.3 80.8 1.7B (154M) - -
@@
Finally, we evaluate zero-shot performance on the OASIS-2 and AIBL datasets. Our pipeline outperforms other fine-tuning strategies and achieves strong zero-shot results without additional fine-tuning. MedBLIP reports high performance on both datasets. However, it should be noted that this method was pre-trained on 27,161 MRIs from ADNI, NACC, and OASIS, making the OASIS evaluation not truly zero-shot.
@@ 4.3.1 Impact of individual component
We assess the effectiveness of each component in our fine-tuning pipeline on ADNI using M3D as the backbone (Table 3). Compared to the baseline (a), fine-tuning the image encoder (b) or text encoder (c) improves accuracy, with the highest achieved when both are fine-tuned (d). Among components, generated reports provide the largest gain, highlighting the benefit of structured metadata, while the MMSE prediction token also contributes to capturing disease-specific patterns. These results demonstrate the complementary contributions of different components.
Given the impact of generated reports in (c), we further examine metadata. Using the fine-tuned model (d), we extract image features from the vision encoder and numeric biomarker features from the text encoder, and train a simple linear classifier on each type and their concatenation to evaluate predictive value independently of the full pipeline. All classifiers are trained and tested on ADNI. As shown in Table 5, biomarkers alone effectively support classification, while combining image and biomarker features further improves performance.
Table 4: Classification accuracy using different input features.
Input features ACC (%)
[…]
Table 5: Classification accuracy under varying fine-tuning sample size on ADNI.
Sample size 500 1000 1500 2000
ACC 63.1% 66.8% 78.8% 81.3%
4.3.2 Impact of fine-tuning sample size
4.3.3 Impact of fine-tuning strategies